elizabeth walks 8 miles for a charity even.She walks half the distance during the first two hours and a fourth of the remaning distance during the next half hour.How far does she have to go to complete the walk?
Answer by mananth(16949)

elizabeth walks 8 miles for a charity even.She walks half the distance during the first two hours and a fourth of the remaning distance during the next half hour.How far does she have to go to complete the walk?
..
2 hours --1/2 the distance = 4 miles.
..
1/2 hour--1/4 of 4 distance = 1 mile.
She has walked 4+1 = 5 miles.
..
So she has to walk 3 more miles.